OpenStudy (anonymous):

Find the area of the surface obtained by rotating the curve y=4x^3 from x=0 to x=4 about the x-axis.

OpenStudy (anonymous):

kinda like a trumpet shape when we rotate the curve right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we can take the radius as being 4x^3

OpenStudy (anonymous):

use the disk method

OpenStudy (anonymous):

\[\int\limits_{0}^{4}\pi[4x^3]^2dx\]

OpenStudy (anonymous):

derivative is 12x^2?

OpenStudy (anonymous):

\[4\pi \int\limits_{0}^{4}x^3dx\]

OpenStudy (anonymous):

find the anti derivative of x^3

OpenStudy (anonymous):

x^4/4?

OpenStudy (anonymous):

yup, which means you got to do: \[\pi x^4\] evaluated from 0 to 4

OpenStudy (anonymous):

that gives me 256pi

OpenStudy (anonymous):

nice job

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

oh this one also needs the surface area, how come you didnt stop me?

OpenStudy (anonymous):

i just realized it.